A 20 ft.lb torque (moment) is applied to a 10 lb drum, which is rotating about its center. The drum has a radius of gyration of 1.5 ft. Determine the magnitude of its angular acceleration.
Answer #1

Torque is given as, T = Iα

k = 1.5ft

I = mk2 = 10 (1.5)2 = 10*2.25 = 22.5 lbft2

T = 20ftlb

20 = 22.5α

or, α = 20/22.5 = 0.888rad/s2